### v1

# SHyPar: A Spectral Coarsening Approach to Hypergraph Partitioning

Abstract
State-of-the-art hypergraph partitioners utilize a multilevel paradigm to construct progressively coarser hypergraphs across multiple layers, guiding cut refinements at each level of the hierarchy. Traditionally, these partitioners employ heuristic methods for coarsening and do not consider the structural features of hypergraphs. In this work, we introduce a multilevel spectral framework, SHyPar, for partitioning large-scale hypergraphs by leveraging hyperedge effective resistances and flow-based community detection techniques. Inspired by the latest theoretical spectral clustering frameworks, such as HyperEF and HyperSF, SHyPar aims to decompose large hypergraphs into multiple subgraphs with few inter-partition hyperedges (cut size). A key component of SHyPar is a flow-based local clustering scheme for hypergraph coarsening, which incorporates a max-flow-based algorithm to produce clusters with substantially improved conductance. Additionally, SHyPar utilizes an effective resistance-based rating function for merging nodes that are strongly connected (coupled). Compared with existing state-of-the-art hypergraph partitioning methods, our extensive experimental results on real-world VLSI designs demonstrate that SHyPar can more effectively partition hypergraphs, achieving state-of-the-art solution quality.
Index Terms:
Hypergraph Partitioning, Effective Resistance, Flow-Based Clustering, Spectral Coarsening.
I Introduction
Hypergraphs are more general than simple graphs since they allow modeling higher-order relationships among the entities. Hypergraph partitioning is an increasingly important problem with applications in various areas, including parallel sparse matrix computations [1], computer-aided design (CAD) of integrated circuit systems [2], physical mapping of chromosomes [3], protein-to-protein interactions [4], as well as data mining and machine learning [5].
The problem of hypergraph partitioning involves grouping the nodes of a hypergraph into multiple clusters. This is done by minimizing a given cost function, defined over the hypergraph, under specific constraints. For instance, the cost function can be the hypergraph cut, which is the number (or total weight) of hyperedges that span more than one partition. The constraints may include balance constraints, where the difference in total node (or hyperedge) weight of each cluster should not exceed a given threshold. Due to these balance constraints, the problem of optimally partitioning a hypergraph is classified as NP-hard [6].
State-of-the-art multilevel hypergraph partitioning techniques predominantly rely on relatively simple heuristics for edge coarsening, such as vertex similarity or hyperedge similarity [2, 7, 8, 9, 10]. For example, hyperedge similarity-based coarsening techniques contract similar large-size hyperedges into smaller ones. While this approach is straightforward to implement, it may adversely affect the original structural properties of the hypergraph. On the other hand, vertex-similarity-based algorithms determine strongly coupled (correlated) node clusters by measuring distances between vertices. This can be facilitated by hypergraph embedding, which maps each vertex to a low-dimensional vector, allowing for the Euclidean distance (coupling) between vertices to be easily computed in constant time. However, these simple metrics often fail to capture the higher-order global (structural) relationships within hypergraphs, potentially leading to suboptimal partitioning solutions.
Spectral methods are increasingly important in various graph and numerical applications [11]. These applications include scientific computing and numerical optimization [12, 13, 14], graph partitioning and data clustering [15, 16, 17], data mining and machine learning [18, 19], graph visualization and data analytics [20, 21, 22], graph signal processing and image segmentation [23, 24, 25], and integrated circuit (IC) modeling and verification [26, 27, 28, 29, 30, 31]. Recent theoretical breakthroughs in spectral graph theory have led to the development of nearly-linear time spectral graph sparsification (edge reduction) [32, 33, 34, 31, 35, 36, 37] and coarsening (node reduction) algorithms [38, 39].
On the other hand, spectral theory for hypergraphs has been less developed due to the more complicated structure of hypergraphs. For example, a mathematically rigorous approach has introduced a nonlinear diffusion process to define the hypergraph Laplacian operator, which measures the flow distribution within each hyperedge [40, 41]; Additionally, the Cheeger’s inequality has been validated for hypergraphs under this diffusion-based nonlinear Laplacian operator [40]. However, these theoretical results do not readily translate into practical implementations. Even recent breakthroughs, such as the SpecPart and K-SpecPart algorithms for spectral hypergraph partitioning [42, 43], still depend heavily on the initial solutions provided by existing multilevel hypergraph partitioning methods, which themselves are based entirely on simplistic edge coarsening heuristics [44, 45].
This paper introduces a brand-new multilevel hypergraph partitioning framework that leverages the latest highly scalable spectral hypergraph coarsening algorithms [46, 47]. Unlike traditional methods that depend solely on simple hyperedge (node) contraction heuristics and focus on local hypergraph structures, our framework, for the first time, incorporates spectral (global) properties into the multilevel coarsening and partitioning tasks as depicted in Figure 1. To achieve these goals, the paper is organized in a structured sequence of steps:
Figure 1: The proposed multilevel hypergraph partitioning via spectral coarsening.
Scalable Spectral Hypergraph Coarsening Algorithms
This paper presents a two-phase scalable algorithmic framework for the spectral coarsening of large-scale hypergraphs, which exploits hyperedge effective resistances and strongly local flow-based methods [46, 47]. The proposed methods facilitate the decomposition of hypergraphs into multiple strongly-connected node clusters with minimal inter-cluster hyperedges, by incorporating the latest diffusion-based nonlinear quadratic operators defined on hypergraphs.
Multilevel Hypergraph Partitioning via Spectral Coarsening
This paper develops a brand-new multilevel hypergraph partitioning tool by seamlessly integrating the proposed spectral coarsening methods into the hypergraph partitioning platform. By replacing the traditional simple coarsening heuristics with our theoretically rigorous spectral methods, the multilevel hypergraph partitioning tools developed through this research will potentially offer significantly improved partitioning solutions without compromising runtime efficiency.
Validations of Multilevel Hypergraph Partitioning Tools
This paper comprehensively validates the developed hypergraph partitioning tools, focusing on an increasingly important application: integrated circuit (IC) computer-aided design. Both the solution quality and runtime efficiency will be carefully assessed by testing the tools on a wide range of public-domain data sets. Additionally, the developed open-source software packages will be made available for public assessment.
The structure of this paper is organized as follows: Section II provides a foundational overview of the essential concepts and preliminaries in spectral hypergraph theory. Section III introduces the proposed method for hypergraph partitioning via spectral clustering, including resistance-based hypergraph clustering, flow-based clustering, and multilevel hypergraph partitioning. Section IV applies our framework to an established hypergraph partitioning tool, showcasing extensive experimental results on various real-world VLSI design benchmarks. The paper concludes with Section V, which summarizes the findings and implications of this work.
II Preliminaries and Background
II-A Spectral (Hyper)graph Theory
II-A1 Graph Laplacian matrix
In an undirected graph $G=(\mathcal{V},\mathcal{E},z)$, the symbol $\mathcal{V}$ represents a set of nodes (vertices), $\mathcal{E}$ represents a set of undirected edges, and $z$ indicates the weights associated with these edges. We denote ${D}$ as a diagonal matrix where each diagonal element ${D}(i,i)$ corresponds to the weighted degree of node $i$. Additionally, ${A}$ is defined as the adjacency matrix for the undirected graph $G$, as described below:
${A}(i,j)=\begin{cases}z(i,j)&\text{ if }(i,j)\in\mathcal{E}\\ 0&\text{otherwise }.\end{cases}$ (1)
Subsequently, the Laplacian matrix of the graph $G$ is determined by the formula $L=D-A$. This matrix adheres to several key properties: (1) the sum of the elements in each row or column is zero; (2) all elements outside the main diagonal are non-positive; (3) the graph Laplacian is symmetric and diagonally dominant (SDD), characterized by non-negative eigenvalues.
II-A2 Courant-Fischer Minimax Theorem
The $k$-th largest eigenvalue of the Laplacian matrix $L\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ over the subspace $U$ of $\mathbb{R}^{\mathcal{V}}$, can be computed as follows:
$\lambda_{k}(L)=\min_{dim(U)=k}\,{\max_{\begin{subarray}{c}x\in U\\ x\neq 0\end{subarray}}{\frac{x^{\top}Lx}{x^{\top}x}}},$ (2)
This can be leveraged for compute the spectrum of the Laplacian matrix $L$.
II-A3 Graph conductance
In a graph $G=(\mathcal{V},\mathcal{E},z)$ where vertices are partitioned into subsets $(S,\hat{S})$, the conductance of partition $S$ is defined as:
$\Phi_{G}(S)\coloneqq\frac{w(S,\hat{S})}{\min\left(vol(S),vol(\hat{S})\right)}=\frac{\sum_{(i,j)\in\mathcal{E}:i\in S,j\notin S}{z(i,j)}}{\min\left(vol(S),vol(\hat{S})\right)},$ (3)
where the volume of the partition, $vol(S)$, is the sum of the weighted degrees of vertices in $S$, defined as $vol(S):=\sum_{i\in S}{d(i)}$. The graph’s conductance [48] is defined as:
$\Phi_{G}:=\min\limits_{\emptyset\not\subseteq S\subseteq\mathcal{V}}\Phi(S).$ (4)
II-A4 Cheegers’ inequality
Research has demonstrated that the conductance $\Phi_{G}$ of the graph $G$ closely correlates with its spectral properties, as articulated by Cheeger’s inequality [48]:
$\omega_{2}/2\leq\Phi_{G}\leq\sqrt{2\omega_{2}},$ (5)
where $\omega_{2}$ is the second smallest eigenvalue of the normalized Laplacian matrix $\widetilde{L}$, defined as $\widetilde{L}=D^{-1/2}LD^{-1/2}$.
II-A5 Effective resistance distance
Let $G=(\mathcal{V},\mathcal{E},z)$ represent a connected, undirected graph with weights $z\in\mathbb{R}^{\mathcal{E}}_{\geq 0}$. let ${b_{p}}\in\mathbb{R}^{\mathcal{V}}$ denote the standard basis vector characterized by zero entries except for a one in the $p$-th position, and let ${b_{pq}}=b_{p}-b_{q}$, The effective resistance between nodes $p$ and $q$, $(p,q)\in\mathcal{V}$ can be computed by:
$R_{eff}(p,q)=b_{pq}^{\top}L_{G}^{\dagger}b_{pq}=\sum\limits_{i=2}^{|\mathcal{V}|}\frac{(u_{i}^{\top}b_{pq})^{2}}{\lambda_{i}}=\max\limits_{x\in\mathbb{R}^{\mathcal{V}}}\frac{(x^{\top}b_{pq})^{2}}{x^{\top}L_{G}x},$ (6)
where $L_{G}^{\dagger}$ represents the Moore-Penrose pseudo-inverse of the graph Laplacian matrix $L_{G}$, and $u_{i}\in\mathbb{R}^{\mathcal{V}}$ for $i=1,...,|\mathcal{V}|$ represents the unit-length, mutually-orthogonal eigenvectors corresponding to Laplacian eigenvalues $\lambda_{i}$ for $i=1,...,|\mathcal{V}|$.
II-A6 Spectral methods for hypergraphs
Classical spectral graph theory shows that the structure of a simple graph is closely related to the graph’s spectral properties. Specifically, Cheeger’s inequality demonstrates the close connection between expansion (or conductance) and the first few eigenvalues of graph Laplacians [16]. Moreover, the Laplacian quadratic form computed with the Fiedler vector (the eigenvector corresponding to the smallest nonzero Laplacian eigenvalue) has been exploited to find the minimum boundary size or cut for graph partitioning tasks [15]. However, there has been very limited progress in developing spectral algorithms for hypergraphs. For instance, a classical spectral method has been proposed for hypergraphs by converting each hyperedge into undirected edges using star or clique expansions [49]. This naive hyperedge conversion scheme may result in lower performance due to ignoring the multi-way high-order relationships between the entities. A more rigorous approach by Tasuku and Yuichi [50] generalized spectral graph sparsification for the hypergraph setting by sampling each hyperedge according to a probability determined based on the ratio of the hyperedge weight to the minimum degree of two vertices inside the hyperedge. Another family of spectral methods for hypergraphs explicitly builds the Laplacian matrix to analyze the spectral properties of hypergraphs. A method has been proposed to create the Laplacian matrix of a hypergraph and generalize graph learning algorithms for hypergraph applications [51]. A more mathematically rigorous approach by Chan et al. introduced a nonlinear diffusion process for defining the hypergraph Laplacian operator by measuring the flow distribution within each hyperedge [40, 41]. Moreover, Cheeger’s inequality has been proven for hypergraphs under the diffusion-based nonlinear Laplacian operator [40].
II-A7 Hypergraph conductance
A hypergraph $H=(V,E,w)$ consists of a vertex set $V$ and a set of hyperedges $E$ with unit weight $w=1$. The degree of a vertex $d_{v}$ is defined as: $d_{v}:=\Sigma_{e\in E:v\in e}w(e)$, where $w(e)$ represents the weight of each hyperedge. The volume of a node set $S\subseteq V$ in the hypergraph is defined as: $vol(S):=\Sigma_{v\in S}d_{v}$. The conductance of a subset $S$ within the hypergraph is then calculated as:
$\Phi(S):=\frac{cut(S,\hat{S})}{min\{vol(S),vol(\hat{S})\}},$ (7)
where $cut(S,\hat{S})$ quantifies the number of hyperedges that cross between $S$ and $\hat{S}$. This computation uses an ”all or nothing” splitting function that uniformly penalizes the splitting of hyperedges. The hypergraph’s overall conductance is defined as:
$\Phi_{H}:=\min\limits_{\emptyset\not\subseteq S\subseteq V}\Phi(S),$ (8)
II-B Hypergraph Partitioning Methods
The previous hypergraph partitioners leverage a multilevel paradigm to construct a hierarchy of coarser hypergraphs using local clustering methods. Computing a sequence of coarser hypergraphs that preserve the structural properties of the original hypergraph is a key step in every partitioning method. The coarsening algorithm in existing partitioning methods either computes matching or clustering at each level by utilizing a rating function to cluster strongly correlated vertices. Hyperedge matching and vertex similarity methods are used in the coarsening phase to cluster the nodes and contract the hyperedges. Existing well-known hypergraph partitioners, such as hMETIS [2], KaHyPar [45], PaToH [9], and Zoltan [7], all use heuristic clustering methods to compute the sequence of coarser hypergraphs.
II-B1 Hypergraph coarsening
Multi-level coarsening techniques typically employ either matchings or clusterings on each level of the coarsening hierarchy. These algorithms utilize various rating functions to decide whether vertices should be matched or grouped together, using the contracted vertices to form the vertex set of the coarser hypergraph at the subsequent level. In contrast, n-level partitioning algorithms, such as the graph partitioner KaSPar [52], establish a hierarchy of (nearly) n levels by contracting just one vertex pair between two levels. This approach eliminates the need for matching or clustering algorithms during the graph reduction process. KaSPar utilizes a priority queue to determine the next vertex pair to be contracted. After each contraction, it updates the priority of every neighboring vertex of the contracted vertex to maintain consistency in priorities. However, in hypergraphs, this method faces significant speed limitations because the size of the neighborhood can greatly expand due to a single large hyperedge. To address this limitation, KaHyPar [45] adopts the heavy-edge rating function. This strategy involves initially selecting a random vertex $p$ and contracting it with the best neighboring node that has the highest rating. The rating function specifically selects a vertex pair ($p,q$) that is involved in a large number of hyperedges with relatively small sizes, optimizing the coarsening process based on the most significant connections between vertices:
$r(p,q)=\sum_{e=(p,q)\in E}{\frac{w(e)}{|e|-1}},$ (9)
where $r(p,q)$ is the rating of the vertex pair, $w(e)$ is the weight of hyperedge $e$, and $|e|$ is the hyperedge cardinality.
II-B2 Community Detection
The coarsening phase aims to generate progressively smaller yet structurally consistent approximations of the input hypergraph. However, certain scenarios may arise where the inherent structure becomes obscured. For example, tie-breaking decisions may be necessary when multiple neighbors of a vertex share the same rating. Consequently, to improve coarsening schemes, existing partitioners like KaHyPar utilize a preprocessing step involving community detection to guide the coarsening phase. In this approach, the hypergraph is divided into several communities and then the coarsening phase is applied to each community separately. Existing community detection algorithms, such as the Louvain algorithm, partition hypergraph vertices into communities characterized by dense internal connections and sparse external ones. This method reformulates the problem into a task of modularity maximization in graphs.
II-B3 Partitioning objectives
Hypergraph partitioning extends the concept of graph partitioning. Its objective is to distribute the vertex set into multiple disjoint subsets while minimizing a specified cut metric and adhering to certain imbalance constraints. The process of dividing into two subsets is known as bipartitioning, whereas dividing into multiple subsets, typically referred to as $k$-way partitioning, involves partitioning into $k$ parts. More formally, consider a hypergraph $H=(V,E,w)$, where $k$ is a positive integer (with $k\geq 2$) and $\epsilon$ is a positive real number (where $\epsilon\leq\frac{1}{k}$). The objective of $k$-way balanced hypergraph partitioning is to divide $V$ into $k$ disjoint subsets $S=\{V_{0},V_{1},…,V_{k-1}\}$ such that:
  • $(\frac{1}{k}-\epsilon)W\leq\sum_{v\in V_{i}}{w_{v}}\leq(\frac{1}{k}+\epsilon)W$, for $10\leq i\leq k-$ 1
  • $cutsize_{H}{(S)}=\sum_{\{e|e\not\subseteq V_{i}\text{ for any }i\}}{w_{e}}$ is minimized
Here, $k$ represents the number of partitions, $W$ is the hypergraph total weight ($W=\sum_{v\in V}{w_{v}}$), $\epsilon$ denotes the allowable imbalance among the partitions, and each $V_{i}$ is a block of the partition. We denote $S$ as an $\epsilon$-balanced partitioning solution.
III SHyPar: Hypergraph Partitioning via Spectral Coarsening
To address the limitations of existing hypergraph coarsening methods that rely on simple heuristics, we proposed a theoretically sound and practically efficient framework for hypergraph coarsening. Specifically, we proposed a two-phase spectral hypergraph coarsening scheme based on the recent research on spectral hypergraph clustering [46, 47]. Phase A utilizes spectral hypergraph coarsening (HyperEF) to decompose a given hypergraph into smaller node partitions with bounded effective-resistance diameters [47]. This is followed by Phase B, which guides the coarsening stage using a flow-based community detection method (HyperSF) aimed at minimizing the ratio cut [46]. Next, we exploit the proposed two-phase spectral hypergraph coarsening method for multilevel hypergraph partitioning: the prior heuristic hypergraph coarsening schemes will be replaced by the proposed spectral coarsening methods to create a hierarchy of coarser hypergraphs that can preserve the key structural properties of the original hypergraph.
III-A Resistance-Based Hypergraph Clustering (Phase A)
The existing coarsening algorithms, contract vertices at each level of the hierarchy. The primary method involves contracting each vertex with the best neighboring node. This is commonly done by using rating functions to identify and contract highly connected vertices. However, these determinations often rely solely on the weights and sizes of the hyperedges. Rating functions, such as those described in Eq. (9), based on hyperedge size are limited to the local structural properties of the hypergraph and do not account for its global structure. In contrast, the effective-resistance diameter provides a more comprehensive criterion, which considers the global structure. Through an example, we show that hyperedge size-based score functions are not optimal, occasionally disrupting the global structure of the hypergraph. For instance, in scenarios where the hypergraph contains a bridge with few nodes (small size hyperedge), as illustrated in Figure 2, algorithms that use hyperedge size tend to inappropriately contract bridge nodes (node 4 and node 7). This contraction can lead to the collapse of the overall structure of the hypergraph. Since the effective resistance of a bridge is high, algorithms based on effective resistance do not contract these nodes and preserve the integrity of the hypergraph structure.
Figure 2: The example of contraction in a simple hypergraph
In [35], the authors introduced a spectral algorithm based on effective resistance to sparsify hypergraphs. This method achieves nearly-linear-sized sparsifiers by sampling hyperedges according to their effective resistances [35]. Despite its theoretical appeal, the technique involves a non-trivial procedure for the estimation of hyperedge effective resistances, which could hinder practical efficiency. The need to convert hypergraphs into graphs through clique expansion and the iterative updating of edge weights significantly adds to the complexity of the algorithm [35].
Figure 3: Overview of the HyperEF method.
III-A1 Hypergraph Clustering by Effective Resistances
A spectral hypergraph coarsening technique, HyperEF, effectively clusters nodes within each hyperedge that have low effective-resistance diameters as shown in Figure 3. This allows for the construction of significantly smaller hypergraphs while preserving the essential structural characteristics of the original hypergraph. A crucial element of HyperEF is an efficient method for estimating hyperedge effective resistances. This is accomplished by adapting the optimization-based effective-resistance estimation method shown in Eq. (6) to hypergraphs. Specifically, the effective resistance of a hyperedge is determined by finding an optimal vector $\chi^{*}$ through a designated optimization process.
$\begin{split}R_{e}(\chi^{*})=\max\limits_{\chi\in\mathbb{R}^{\mathcal{V}}}\frac{(\chi^{\top}b_{pq})^{2}}{Q_{H}(\chi)},\quad p,q\in e\end{split}$ (10)
where $x^{\top}L_{G}x$ in Eq. (6) is replaced by the following nonlinear quadratic form $Q_{H}(\chi)$ [40]:
$Q_{H}(\chi):=\sum\limits_{e\in E}w_{e}\max\limits_{u,v\in e}(\chi_{u}-\chi_{v})^{2}.$ (11)
HyperEF generates a significantly smaller hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ from the original hypergraph $H=(V,E,w)$ by utilizing hyperedge effective resistances, achieving reductions in the numbers of vertices, edges, and weights ($|V^{\prime}|<|V|$, $|E^{\prime}|<|E|$, and $|w^{\prime}|<|w|$).
III-A2 Low-Resistance-Diameter Hypergraph Decomposition
Consider a weighted undirected graph $G=(\mathcal{V},\mathcal{E},z)$ where the weights $z\in\mathbb{R}^{\mathcal{E}}{>0}$ and $\gamma>1$ is sufficiently large. The effective-resistance diameter is defined as $\max\limits_{u,v\in\mathcal{V}}R_{eff}(u,v)$. Recent studies demonstrate that it is feasible to partition a simple graph $G$ into several node clusters $G[\mathcal{V}_{i}]$ that each have a low effective-resistance diameter, by removing only a constant fraction of the edges [53]:
$\max\limits_{u,v\in\mathcal{V}_{i}}R_{eff_{G[\mathcal{V}_{i}]}}(u,v)\lesssim\gamma^{3}\frac{|\mathcal{V}|}{z(\mathcal{E})}.$ (12)
Consider the weighted undirected graph $G=(\mathcal{V},\mathcal{E},z)$ where the weights $z\in\mathbb{R}^{\mathcal{E}}_{\geq 0}$, and let $\Phi_{G}$ represent the conductance of $G$. Cheeger’s inequality provides a way to establish a relationship between the effective-resistance diameter of the graph and its conductance [53]:
$\max\limits_{u,v\in\mathcal{V}}R_{eff}(u,v)\lesssim\frac{1}{\Phi_{G}^{2}}.$ (13)
The HyperEF algorithm is based on recent spectral hypergraph theory [40, 41] by extending the above theorem to the hypergraph setting: inequality (12) implies that it is possible to decompose a hypergraph into multiple (hyperedge) clusters that have small effective-resistance diameters by removing only a few inter-cluster hyperedges, while (13) implies that contracting the hyperedges (node clusters) with small effective-resistance diameters is not significantly impacting the original hypergraph conductance. HyperEF allows computing a much smaller hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ given the original hypergraph $H=(V,E,w)$ by exploiting hyperedge effective resistances, which consists of the following three key steps:
Step 1
constructs the Krylov subspace to approximate the eigensubspace related to the original hypergraph;
Step 2
estimates the effective resistance of each hyperedge by applying the proposed optimization-based method;
Step 3
constructs the coarsened hypergraph by aggregating node clusters with low effective-resistance diameters.
III-A3 Fast Estimation of Hyperedge Effective Resistances
To enhance efficiency, the search for $\chi^{*}$ in Eq. (10) is restricted to an eigensubspace represented by a few select Laplacian eigenvectors from the simplified graph derived from the original hypergraph. Consider $G_{b}=(\mathcal{V}_{b},\mathcal{E}_{b},z_{b})$ serves as the bipartite graph equivalent to the hypergraph $H=(V,E,w)$, where $|\mathcal{V}_{b}|=|V|+|E|$, $|\mathcal{E}_{b}|=\Sigma_{e\in E}|e|$, and $z_{b}$ denotes the scaled edge weights, specifically: $z(e,p)=\frac{w(e)}{d(e)}$. According to Eq. (6) computing approximate hyperedge effective resistances involves identifying a set of orthogonal eigenvectors that optimize (10).
Figure 4: The approximated effective resistances for a simple graph.
To avoid the computational complexity of determining eigenvalues and eigenvectors, we have used a scalable algorithm to approximate the eigenvectors by leveraging the Krylov subspace, defined as follows:
For a non-singular matrix $A_{n\times n}$, and a non-zero vector $\chi\neq 0\in\mathbb{R}^{n}$, the order-$(\rho+1)$ Krylov subspace generated by $A$ from $x$ is
$\kappa_{\rho}(A,x):=span(x,Ax,A^{2}x,...,A^{\rho}x),$ (14)
where $x$ is a random vector, and $A$ signifies the normalized adjacency matrix of the simple graph translated from the hypergraph via star expansion. Note that for each hypergraph, the Krylov subspace vectors in Eq. (14) only need to be computed once, which can be achieved in nearly linear time using only sparse matrix-vector operations. The effective resistance of each hyperedge can be estimated in constant time by identifying a few Krylov subspace vectors that maximize Eq. (10). As shown in Figure 4, the approximate effective resistances of simple graphs obtained using a few Krylov vectors can well correlate with the ground truths. Algorithm 1 provides the detailed flow of the proposed hyperedge effective resistance estimation method.
Algorithm 1 The effective resistance estimation algorithm flow
Input: Hypergraph $H=(V,E,w)$, $\rho$.
Output: A vector of effective resistance $R$ with the size $|E|$.
1:  Construct the bipartite graph $G_{b}$ corresponding to $H$.
2:  Construct the order-$(\rho+1)$ Krylov subspace.
3:  Use Gram–Schmidt method to obtain the orthogonal vectors.
4:  For each hyperedge compute its $\rho$ resistance ratios using (15).
5:  Obtain all hyperedge effective resistances $R$ based on (17).
6:  Return $R$.
It is assumed that $x^{(1)},x^{(2)},...,x^{(\rho)}\in\mathbb{R}^{\mathcal{V}_{b}}$ are the $\rho$ mutually-orthogonal vectors based on the order-$(\rho+1)$ Krylov subspace $\kappa_{\rho}(A,x)$, which is constructed in Step 1. The effective resistance estimation method, as described in (6), is extended by the HyperEF algorithm through the incorporation of the non-linear quadratic operator of hypergraphs (11), to include hypergraph spectral properties. By excluding the node embedding values associated with the star nodes from $x^{(i)}$, a new set of vectors $\chi^{(i)}$, all mutually orthogonal, is generated. Subsequently, each node in the hypergraph is embedded into a $\rho$-dimensional space. The resistance ratio ($r_{e}$) associated with a vector $\chi^{(i)}\in\chi^{(1)},...,\chi^{(\rho)}$ for each hyperedge $e$ is then computed as follows:
$r_{e}(\chi^{(i)})=\frac{(\chi^{(i)\top}b_{pq})^{2}}{Q_{H}(\chi^{(i)})},\quad p,q\in e$ (15)
where $p$ and $q$ are identified as the two maximally-separated nodes in the $\rho$-dimensional embedding space. Multiple resistance ratios $r_{e}^{(1)},...,r_{e}^{(\rho)}$ are returned by Eq. (10), corresponding to $\chi^{(1)},...,\chi^{(\rho)}$.
After the resistance ratios are sorted in descending order, it is observed that:
$r_{e}^{1}>r_{e}^{2}>...>r_{e}^{\rho}.$ (16)
In HyperEF, the top $m$ resistance ratios are selected to estimate the effective resistance of each hyperedge. The hyperedge effective resistance ($R_{e}$) is specifically approximated by:
$R_{e}=\sum\limits_{i=1}^{m}r_{e}^{i},\quad e\in E.$ (17)
III-A4 Practical Implementations
To achieve good efficiency, the proposed spectral hypergraph coarsening framework leverages a linear-time local spectral embedding scheme based on low-pass filtering (smoothing) of random graph signals [54, 55, 25, 39, 56]. The key challenge of the above framework lies in designing universal low-pass filtering functions applicable to hypergraphs. To this end, low-pass graph filters for hypergraphs are designed based on the latest research in spectral hypergraph theory [57, 58, 50, 40]. One promising research direction is to convert hypergraphs into simple undirected graphs so that existing graph filters can be directly adopted for constructing the Krylov subspace required by the HyperEF algorithm. In addition, a spectrum-preserving edge expansion scheme is investigated based on: 1) clique expansion (CE), which replaces each hyperedge with multiple edges forming a clique among the incident vertices of the hyperedges; 2) star expansion (SE), which replaces each hyperedge with a new vertex connected to its incident vertices; 3) hybrid expansion (HE), which adaptively applies CE or SE. Specifically, the proposed spectral method determines whether a hyperedge should be replaced by a star or clique expansion according to the nonlinear diffusion operators defined on the hypergraph [40, 41], to achieve an optimal balance between complexity and quality. Compared with recent theoretical results in spectral methods for hypergraphs [50] that may not be efficiently implemented due to the polynomial-time algorithm complexity, the proposed spectral coarsening algorithms can be implemented using parallel-friendly sparse kernels, which is suitable for accelerations on modern hardware platforms, such as CPUs, GPUs, and FPGAs [59, 60, 61, 62, 63, 64, 65, 66].
III-A5 Multilevel Effective Resistance Clustering
We enhance the hypergraph structure by employing a technique that iteratively computes a vector of effective resistance $R$ and contracts hyperedges exhibiting low effective resistances ($R_{e}<\delta$), where $\delta$ is the effective resistance threshold. Throughout this process, node clusters are contracted by merging nodes within each cluster and replacing them with a new entity termed a supernode at the subsequent level. The structural information of the hypergraph is preserved and transmitted through levels by assigning weights to the supernodes, which are equal to the hyperedge effective resistances evaluated at the preceding level.
Let $H^{(l)}=(V^{(l)},E^{(l)},w^{(l)})$ represent the hypergraph at the $l$-th level, where $S\in V^{(l)}$ is a cluster of nodes producing a supernode $\vartheta\in V^{(l+1)}$ at level $l+1$. The vector of node weights is denoted as $\eta^{(l)}\in\mathbb{R}_{\geq 0}^{V^{(l)}}$, initially set to all zeros for the original hypergraph. At each level $l$, $\eta$ is updated by:
$\eta_{\vartheta}:=\sum\limits_{j=1}^{|S|}\eta(v_{j}^{(l)}).\quad v\in V^{(l)}:v\in S$ (18)
III-A6 Node Weight Propagation (NWP)
The vector of effective resistance $R$ is updated based on the node weights $\eta$ to transmit the clustering information from previous levels to the current level:
$R_{e}^{(l)}=\sum\limits_{k=1}^{|e|}\eta(v_{k}^{(l)})+R_{e}^{(l)}.$ (19)
Consequently, the effective resistance of a hyperedge at a coarser level depends not only on the effective resistance ($R_{e}^{(l)}$) evaluated at the current level, but also on the data transferred from all previous (finer) levels.
The experimental results indicate that using Eq. (19) for effective resistances estimation yields more balanced hypergraph clustering outcomes compared to approaches that ignores the previous clustering information. The complete workflow for the effective resistance clustering algorithm, HyperEF, used for coarsening a hypergraph $H$ across $L$ levels, is detailed in Algorithm 2.
Algorithm 2 The HyperEF algorithm for hypergraph clustering flow
Input: Hypergraph $H=(V,E,w)$, $\delta$, $L$, $\eta$.
Output: A coarsened hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ that $|V^{\prime}|\ll|V|$.
1:  Initialize $H^{\prime}$leftarrow $H$
2:  for $l\leftarrow 1$ to $L$ do
3:    Call Algorithm 1 to compute a vector of effective resistance $R$ with the size $|E^{\prime}|$ for given hypergraph $H^{\prime}$.
4:    Compute the node weights using (18).
5:    Update the effective resistance vector $R$ by applying (19).
6:    Sort the hyperedges with ascending $R$ values.
7:    Starting with the hyperedges that have the lowest effective resistances, contract (cluster) the hyperedge (nodes) if $R_{e}<\delta$.
8:    Construct a coarsened hypergraph $H^{\prime}$ accordingly.
9:  end for
10:  Return $H^{\prime}$.
III-B Flow-Based Community Detection (Phase B)
To improve coarsening schemes for hypergraph partitioning, we utilize a community structure that integrates global hypergraph information into the coarsening process. This community structure directs the coarsening phase by permitting contractions solely within clusters.
In the flow-based community detection, we employ multilevel clustering through HyperEF, which enhances the spectral hypergraph clustering method by integrating a multilevel coarsening approach. Let $H=(V,E,w)$, the hypergraph local conductance (HLC) with respective to a node-set $S$ is defined as follows [67]:
$\textbf{HLC}_{\mathcal{C}}(S)=\frac{cut(S,\hat{S})}{vol(S\cap\mathcal{C})-\beta vol(S\cap\hat{\mathcal{C}})},$ (20)
where $\mathcal{C}\subseteq V$ is reference node set, and $\beta$ is a locality parameter that modulates the penalty for incorporating nearby nodes outside set $\mathcal{C}$.
A spectral hypergraph coarsening algorithm (HyperSF) is proposed by minimizing the HLC, which has achieved promising results in hypergraph coarsening and partitioning in realistic VLSI designs.
Figure 5: Overview of the HyperSF method.
III-B1 Overview of Coarsening Refinement (HyperSF)
Figure 5 shows an overview of the HyperSF method . In this work, HyperSF is leveraged for only refining the most imbalanced node clusters (initially identified by HyperEF) with significantly smaller resistance diameters compared to the rest. Specifically, HyperSF aggregates strongly-coupled node clusters via minimizing Eq. (20): for each selected node set with large imbalance, HyperSF repeatedly solves a max $s$-$t$ flow, min $s$-$t$ cut problem to detect a set of neighboring node clusters that minimizes the local conductance HLC in Eq. (20). To this end, the following key steps are applied (as shown in Figure 6): (Step 1) an auxiliary hypergraph is constructed by introducing a source vertex $s$ and sink vertex $t$; (Step 2) each hyperedge is replaced with a directed graph; (Step 3) each seed node-set is iteratively updated by including new nodes into the set to minimize the HLC by repeatedly solving the max $s$-$t$ flow, min $s$-$t$ cut problem:
$cut^{s-t}(S)=cut_{H}(S)+vol_{H}(\hat{S}\cap\mathcal{C})+\beta vol_{H}(S\cap\hat{\mathcal{C}}).$ (21)
(Step 4) The node sets obtained from flow-based methods that minimize the local conductance are exploited to produce a smaller hypergraph with fewer nodes while preserving the key structural properties of the original hypergraph.
Figure 6: A hyperedge (top) and a hypergraph (bottom) converted to the corresponding directed graphs.
Figure 7: HyperSF minimizes HLC by iteratively solving local max-flows.
III-B2 Local clustering algorithms
An algorithm is local if the input is a small portion of the original dataset. The HyperSF algorithm can be made strongly local when expanding the network around the seed nodes $\mathcal{C}$, which obviously benefits the proposed hypergraph coarsening framework: (1) applying the max $s$-$t$ flow, min $s$-$t$ cut problem on the local neighborhood of the seed nodes restricts node-aggregation locally and keeps the global hypergraph structure intact; (2) such a local clustering scheme significantly improves the algorithm efficiency due to the small-scale input dataset.
III-B3 Flow-based Local Clustering in HyperSF
First, we apply HyperEF (Algorithm 2) to the hypergraph $H=(V,E,w)$ to compute a coarsened hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ and identify isolated nodes, denoted by $\mathcal{C}$. To achieve flow-based local clustering of hypergraph nodes, HyperSF then constructs a sub-hypergraph $H^{\prime}_{L}$ by iteratively expanding the hypergraph around the seed node set $\mathcal{C}$ and then repeatedly solve the hypergraph cut problem to minimize HLC until no significant changes in local conductance are observed. Define $E^{\prime}(S)=\cup_{v^{\prime}\in V^{\prime},v^{\prime}\in\{e^{\prime}\}_{e^{\prime}\in E^{\prime}}}E^{\prime}(v)$ for any set $S\subseteq V^{\prime}$ and let $H^{\prime}_{\varsigma}=(V^{\prime}\cup\{s,t\},E^{\prime}\cup E^{\prime st})$, where $E^{\prime st}$ denotes the terminal edge set. HyperSF aims to construct a sub-hypergraph $H^{\prime}_{L}$ to replace $H^{\prime}_{\varsigma}$ that minimizes HLC by repeatedly solving a local version of (21). To this end, HyperSF sets up an oracle to discover a set of best neighborhood vertices for a given vertex $v^{\prime}$:
$\kappa(v^{\prime})=\{u^{\prime}\in V^{\prime}\}_{(u^{\prime},v^{\prime})\in e^{\prime},e^{\prime}\in E^{\prime}}.$ (22)
HyperSF lets the oracle accept a set of seed nodes $\mathcal{C}$ and return $\kappa(\mathcal{C})=\cup_{v^{\prime}\in\mathcal{C}}\kappa(v^{\prime})$. By utilizing the best neighborhood of the seed nodes $\kappa(\mathcal{C})$, HyperSF builds a local hypergraph $H^{\prime}_{L}=(V^{\prime}_{L}\cup\{s,t\},E^{\prime}_{L}\cup{E^{\prime}}_{L}^{st})$, where $V^{\prime}_{L}=\mathcal{C}\cup\kappa(\mathcal{C})$ and $E^{\prime}_{L}=\{e^{\prime}\in E^{\prime}\mid V^{\prime}_{L}\in e^{\prime}\}$. As shown in Figure 7, HyperSF creates the local auxiliary hypergraph of $H^{\prime}_{L}$ by introducing the source node $s$ and the sink node $t$, so that ${E^{\prime}}_{L}^{st}\subseteq{E^{\prime}}^{st}$ and repeatedly solves the max $s$-$t$ flow, min $s$-$t$ cut problem to minimize HLC. The algorithm continuously expands $H^{\prime}_{L}$ and includes more vertices and hyperedges from $H^{\prime}_{\varsigma}$ by solving (21) for the local hypergraph $H^{\prime}_{L}$.
Algorithm 3 Flow-based Hypergraph Clustering
Input: Hypergraph $H=(V,E,w)$, and $\xi$ (Convergence parameter)
Output: A set of vertices $S$ that minimizes HLC(S);
1:  Apply HyperEF to $H$ to compute coarsened hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$, and isolated nodes $\mathcal{C}\subseteq V$;
2:  Assign isolated nodes as seed nodes $S\leftarrow\mathcal{C}$;
3:  $\Delta_{\textbf{HLC}}\leftarrow\infty$;
4:  while $\Delta_{\textbf{HLC}}>\xi$ do
5:    Identify the best neighborhood of seed nodes $\kappa(S)$;
6:    Update $S$ according to $\kappa(S)$ to construct $H^{\prime}_{L}$;
7:    Add a source node $s$ and sink node $t$ to $H^{\prime}_{L}$;
8:    Repeatedly solve the max $s$-$t$ flow, min $s$-$t$ cut problem by minimizing (21) for $H^{\prime}_{L}$;
9:    $\Delta_{\textbf{HLC}}\leftarrow\textbf{HLC(S)}^{j}-\textbf{HLC(S)}^{j-1}$;
10:  end while
11:  Return $S$.
The algorithm 3 presents the details of the flow-based local clustering technique. It accepts the original hypergraph $H=(V,E,w)$, and the convergence parameter $\xi$, which will output a set of strongly connected vertices $S$ that minimizes HLC.
III-C Algorithm Complexity for Spectral Coarsening
In HyperEF, the complexity for constructing the Krylov subspace for the bipartite graph $G_{b}=(\mathcal{V}_{b},\mathcal{E}_{b},z_{b})$ corresponding to the original hypergraph $H=(V,E,w)$ is $O(|\mathcal{E}_{b}|)$; the complexity of the hyperedge effective resistance estimation and hyperedge clustering is $O(\rho|E|)$; the complexity of computing the node weights through the multilevel framework is $O(|E|)$ that leads to the overall nearly-linear algorithm complexity of $O(\rho|E|+|\mathcal{E}_{b}|)$. In HyperSF, the runtime complexity of the proposed strongly local flow-based algorithm is $O\left(k^{3}vol_{H}(\mathcal{C})^{3}(1+\epsilon^{-1})^{3}\right)$, where $k$ is the maximum hyperedge cardinality. Since each phase in the proposed spectral hypergraph coarsening method has a nearly linear-time complexity, the entire two-phase spectral coarsening procedure will be highly scalable for handling large-scale hypergraphs.
III-D Hypergraph Partitioning with Spectral Coarsening
III-D1 Multilevel spectral coarsening
The spectral hypergraph coarsening method constructs a sequence of successively smaller (coarser) but spectrally-similar hypergraphs via repeatedly contracting groups of vertices (clusters). Initially, the node spectral embeddings are computed by applying the Krylov subspace method in linear time to obtain the vectors of various spectra. The optimization-based effective-resistance estimation formulation is extended to the hypergraph settings by leveraging the nonlinear quadratic operator of the hypergraphs. Then, a low-resistance-diameter hypergraph decomposition method determines the highly connected vertices in multi-dimensional spectral space.
We introduce a new rating function based on the effective resistance formula, which selects a pair of vertex ($p,q$) from hyperedges that have a large number of heavy nets with low effective resistance.
$r(p,q)=\sum_{e=(p,q)\in E}{\frac{w(e)}{|R_{e}|-1}},$ (23)
where $r(p,q)$ is the rating of the vertex pair, $w(e)$ is the weight of hyperedge $e$, and $|R_{e}|$ is the hyperedge effective resistance calculated by Eq. (17).
III-D2 Multilevel coarsening with HyperEF
The resistance-based hypergraph clustering algorithm allows transferring hypergraph structural information through the levels by assigning a weight (that is equal to the effective-resistance diameter of each cluster evaluated at the previous level) to the new corresponding vertices. As a result, the effective hyperedge resistance at a coarse level depends on not only the evaluated effective resistance computed at the current level but also the results transferred from all the previous (finer) levels.
Figure 8: Multilevel spectral coarsening.
As an example, Figure 8 illustrates the two-level coarsening of a hypergraph with 9 vertices and 4 hyperedges. Initially, the vertices are clustered into 4 clusters $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, and the vertices within each cluster are aggregated to create a coarser hypergraph with 4 vertices and 3 hyperedges. In the next level, 4 vertices are grouped into 2 clusters $C^{\prime}_{1}$ and $C^{\prime}_{2}$ that form a coarser hypergraph with 2 vertices and 1 hyperedge.
III-D3 Hypergraph clustering with HyperSF
The flow-based clustering algorithm is a clustering refinement technique designed to enhance the clusters generated by HyperEF. It sorts the clusters based on their resistance diameters, starting with those that have low resistance diameters. HyperSF then creates new clusters by merging those with low-resistance diameters within the same neighborhood, utilizing a flow-based clustering technique. It is important to note that clusters with low resistance diameters typically indicate a strong connection among the nodes within the cluster, and therefore, merging these nodes is unlikely to significantly alter the global structure of the hypergraph. Ultimately, HyperSF aims to produce clusters characterized by low hypergraph local conductance (HLC) to preserve the spectral properties of the original hypergraph.
In our approach, we perform community detection on hypergraphs by translating this challenge into a hypergraph clustering problem. Our algorithm identifies the node cluster with the smallest effective-resistance diameter and refines it by applying a strongly-local flow-based method (HyperSF). This is followed by a community-aware coarsening phase that applies the hypergraph coarsening algorithm to each community separately.
IV Experimental Validation
There are many applications related to hypergraph partitioning. This paper focuses on comprehensively evaluating the performance of the proposed hypergraph partitioning framework for increasingly important applications related to integrated circuits computer-aided design. Both the solution quality and runtime efficiency will be carefully assessed by testing them on a wide range of public-domain data sets. The developed open-source software packages will also be made available for public assessment.
To assess the performance of the proposed multilevel hypergraph partitioning tools in applications related to VLSI designs, We apply the multilevel hypergraph partitioning tool developed to partition public-domain VLSI design benchmarks. For example, the ISPD98 benchmarks that include “IBM01”, “IBM02”, …, “IBM18” hypergraph models with $13,000$ to $210,000$ nodes be adopted [68]. The performance metrics for multilevel hypergraph partitioning, including the total hyperedge cut, imbalance factors, and runtime efficiency, will be considered for comparisons with state-of-the-art hypergraph partitioning tools, such as hMETIS [2], and KaHyPar [45].
IV-A Spectral coarsening with HyperEF and HyperSF
In this section, we compare the preliminary implementations of resistance-based hypergraph clustering (HyperEF) and flow-based clustering (HyperSF) with the well-known hypergraph partitioning tool, hMETIS [2]. Real-world VLSI design (hypergraph) benchmarks have been tested [68]. All experiments have been evaluated on a computing platform with 8 GB of RAM and a 2.2 GHz Quad-Core Intel Core i7 processor.
Figure 9: Resistance-based vs hMETIS Clustering Results.
IV-A1 HyperEF vs hMETIS for Hypergraph Coarsening
HyperEF is compared to hMETIS for hypergraph coarsening considering both solution quality and runtime efficiency. The following average conductance of the node clusters is used to analyze the performance of each method.
$\Phi_{\text{avg}}=\frac{1}{\left|S\right|}\sum\limits_{i=1}^{\left|S\right|}\Phi(S_{i})$ (24)
Where $\Phi(S_{i})$ denotes the conductance of node cluster $S_{i}$. Figure 9 demonstrates the node clustering results for a small hypergraph obtained using HyperEF and hMETIS. Both methods partition the hypergraph into four clusters, and the average conductance of node clusters has been computed to evaluate the performance of each method. The results show that HyperEF outperforms hMETIS by creating node clusters with a significantly lower average conductance. In addition, Table I shows the average conductance of node clusters $\Phi_{\text{avg}}$ computed with both HyperEF and hMETIS by decomposing the hypergraph with the same node reduction ratios (NRs). With an NR = $75\%$ (3×3$\times$3 × node reduction) HyperEF outperforms hMETIS in average conductance while achieving 2438×24-38$\times$24 - 38 × speedups over hMETIS.
TABLE I: HyperEF vs hMETIS coductance (NR=75%)
Benchmark
$\Phi_{\text{avg}}$
HyperEF
$\Phi_{\text{avg}}$
hMETIS
$\mathcal{T}$(seconds)
HyperEF
$\mathcal{T}$(seconds)
hMETIS
IBM01 0.62 0.65 1.23 29 (24$\times$)
IBM02 0.62 0.67 1.41 49 (35$\times$)
IBM03 0.63 0.66 2.11 53 (25$\times$)
IBM04 0.64 0.66 2.37 60 (25$\times$)
IBM05 0.59 0.63 2.34 62 (26$\times$)
IBM06 0.64 0.66 2.63 78 (30$\times$)
IBM07 0.63 0.67 3.54 115 (32$\times$)
IBM08 0.61 0.67 4.15 125 (30$\times$)
IBM09 0.64 0.66 4.38 131 (30$\times$)
IBM10 0.63 0.67 5.79 181 (31$\times$)
IBM11 0.64 0.67 5.73 176 (31$\times$)
IBM12 0.65 0.7 5.94 191 (32$\times$)
IBM13 0.65 0.68 6.87 229 (33$\times$)
IBM14 0.62 0.66 11.51 393 (34$\times$)
IBM15 0.66 0.69 14.44 486 (34$\times$)
IBM16 0.63 0.67 14.62 533 (36$\times$)
IBM17 0.66 0.7 15.22 568 (37$\times$)
IBM18 0.6 0.67 15.79 602 (38$\times$)
IV-A2 HyperSF vs hMETIS for Hypergraph Coarsening
In this section, we evaluate the performance of HyperSF against the hMETIS hypergraph partitioning tool. We measure the average local conductance $\text{{HLC}}_{\text{avg}}$ of the node clusters generated by each method, calculated as follows:
$\text{{HLC}}_{\text{avg}}=\frac{1}{\left|S\right|}\sum\limits_{i=1}^{\left|S\right|}\textbf{HLC}(S^{i}).$ (25)
Table II presents the average local conductance $\text{{HLC}}_{\text{avg}}$ for various methods under the same hypergraph reduction ratio (RR), where we reduce the number of nodes in each original hypergraph by $75\%$. The experimental data illustrate that HyperSF significantly enhances the average local conductance compared to the hMETIS method in all test scenarios.
TABLE II: HyperSF vs hMETIS local coductance (NR=75%)
Benchmark
$\text{{HLC}}_{\text{avg}}$
HyperSF
$\text{{HLC}}_{\text{avg}}$
hMETIS
$\mathcal{T}$(seconds)
HyperSF
$\mathcal{T}$(seconds)
hMETIS
IBM01 0.44 0.65 9.4 29 (3$\times$)
IBM02 0.52 0.69 22.6 49 (2$\times$)
IBM03 0.48 0.67 14.1 53 (4$\times$)
IBM04 0.47 0.68 15 60 (4$\times$)
IBM05 0.55 0.65 29.2 62 (2$\times$)
IBM06 0.51 0.68 30.1 78 (3$\times$)
IBM07 0.48 0.68 26.4 115 (4$\times$)
IBM08 0.48 0.68 43.3 125 (3$\times$)
IBM09 0.47 0.69 24.5 131 (5$\times$)
IBM10 0.48 0.68 45.2 181 (4$\times$)
IBM11 0.46 0.69 30.1 176 (6$\times$)
IBM12 0.50 0.71 42.4 191 (5$\times$)
IBM13 0.48 0.69 50.4 229 (5$\times$)
IBM14 0.48 0.67 85.7 393 (5$\times$)
IBM15 0.47 0.71 96 486 (5$\times$)
IBM16 0.50 0.70 116.8 533 (5$\times$)
IBM17 0.51 0.73 141.3 568 (4$\times$)
IBM18 0.46 0.68 129 602 (5$\times$)
IV-B Hypergraph Partitioning with Spectral Coarsening
We compared SHyPar with leading hypergraph partitioners hMETIS [2], SpecPart [43], KaHyPar [45], and MedPart [69] using two sets of publicly available benchmarks: the ISPD98 VLSI Circuit Benchmark Suite [68] and the Titan23 Suite [70]. The details of these benchmarks are outlined in Table III and Table IV. All tests were conducted on a server equipped with Intel(R) Xeon(R) Gold 6244 processors and an NVIDIA Tesla V100S GPU with 1546GB of memory.
IV-B1 Experimental Setup
To implement SHyPar, we have developed new hypergraph partitioning tools based on the existing open-source multilevel hypergraph partitioner, KaHyPar. We are utilizing the proposed two-phase spectral hypergraph coarsening method. Specifically, the heuristic coarsening scheme has been replaced by our novel spectral coarsening algorithm to create a hierarchy of coarser hypergraphs that preserve the key structural properties of the original hypergraph. Accordingly, we have substituted the existing coarsening method in KaHyPar with our proposed method, incorporating a new rating function. Additionally, the existing algorithm for community detection, Louvain, has been replaced with our proposed flow-based community detection method (Phase 2).
IV-B2 SHyPar Performance on ISPD98 Benchmarks
Table III presents a comparison of the cut sizes achieved by SHyPar on the ISPD98 VLSI circuit benchmark against those obtained from hMETIS, SpecPart, KaHyPar, and MedPart. The results for SHyPar show an average improvement of approximately 0.54% for $\epsilon$ = $2\%$ and 0.4% for $\epsilon$ = $10\%$, affirming its superiority over the best-published results. In several instances, SHyPar outperforms the best-published results by up to 5%; these instances are specifically underlined for emphasis. Figure 10 depicts the cut sizes obtained by SHyPar, KaHyPar, and hMETIS, normalized against the KaHyPar results. It is evident that SHyPar significantly enhances performance over both KaHyPar and hMETIS across many tests. Moreover, when SHyPar was applied to four partitions with $\epsilon$ = $1\%$, the improvements were consistent, as demonstrated in Figure 11, which compares the cut sizes with those from KaHyPar and hMETIS, also normalized by KaHyPar results.
TABLE III: Statistics of ISPD98 VLSI circuit benchmark suite. The best results among all the methods are colored red.
Benchmark Statistics $\epsilon$ = $2\%$ $\epsilon$ = $10\%$
$|V|$ $|E|$ SpecPart hMETIS KaHyPar MedPart SHyPar SpecPart hMETIS KaHyPar MedPart SHyPar
IBM01 12,752 14,111 202 213 202 202 201 171 190 173 166 166
IBM02 19,601 19,584 336 339 328 352 327 262 262 262 264 262
IBM03 23,136 27,401 959 972 958 955 952 952 960 950 955 950
IBM04 27,507 31,970 593 617 579 583 579 388 388 388 389 388
IBM05 29,347 28,446 1720 1744 1712 1748 1707 1688 1733 1645 1675 1645
IBM06 32,498 34,826 963 1037 963 1000 969 733 760 735 788 733
IBM07 45,926 48,117 935 975 894 913 882 760 796 760 773 760
IBM08 51,309 50,513 1146 1146 1157 1158 1140 1140 1145 1120 1131 1120
IBM09 53,395 60,902 620 637 620 625 620 519 535 519 520 519
IBM10 69,429 75,196 1318 1313 1318 1327 1254 1261 1284 1250 1259 1244
IBM11 70,558 81,454 1062 1114 1062 1069 1051 764 782 769 774 763
IBM12 71,076 77,240 1920 1982 2163 1955 1986 1842 1940 1841 1914 1841
IBM13 84,199 99,666 848 871 848 850 831 693 721 693 697 655
IBM14 147,605 152,772 1859 1967 1849 1876 1842 1768 1665 1534 1639 1534
IBM15 161,570 186,608 2741 2886 2737 2896 2728 2235 2262 2135 2169 2135
IBM16 183,484 190,048 1915 2095 1952 1972 1887 1619 1708 1619 1645 1619
IBM17 185,495 189,581 2354 2520 2284 2336 2285 1989 2300 1989 2024 1989
IBM18 210,613 201,920 1535 1587 1915 1955 1521 1537 1550 1915 1829 1520
IBM01IBM02IBM03IBM04IBM05IBM06IBM07IBM08IBM09IBM10IBM11IBM12IBM13IBM14IBM15IBM16IBM17IBM18707070707575757580808080858585859090909095959595100100100100105105105105110110110110Normalized Cutsize (%)KaHyParSHyParhMETIS
Figure 10: ISPD98 benchmarks with unit weights $\epsilon$ = $2\%$ k = 2.
IBM01IBM02IBM03IBM04IBM05IBM06IBM07IBM08IBM09IBM10IBM11IBM12IBM13IBM14IBM15IBM16IBM17IBM18707070707575757580808080858585859090909095959595100100100100105105105105110110110110Normalized Cutsize (%)KaHyParSHyParhMETIS
Figure 11: ISPD98 benchmarks with unit weights $\epsilon$ = $1\%$ k = 4.
IV-B3 SHyPar Performance on Titan23 Benchmarks
Building on the findings from the ISPD98 benchmarks, our research extended to the Titan23 benchmarks, notable for their high-degree hyperedges. Table IV details the results on the Titan23 benchmarks, where SHyPar dramatically outperforms hMETIS, achieving a 12% improvement for $\epsilon$ = $2\%$ and an impressive 25% for $\epsilon$ = $20\%$. Also SHyPar outperforms SpecPart, achieving a 3% improvement for $\epsilon$ = $2\%$ and 1% for $\epsilon$ = $20\%$. In cases such as sparcT2_core, SHyPar even exceeds the best published results by up to 15%; these notable achievements are highlighted by underlining.
TABLE IV: Statistics of Titan23 benchmark suite and cut sizes of different approaches. The best results among all the methods are colored red.
Benchmark Statistics $\epsilon$ = $2\%$ $\epsilon$ = $20\%$
$|V|$ $|E|$ SpecPart hMETIS KaHyPar MedPart SHyPar SpecPart hMETIS KaHyPar MedPart SHyPar
sparcT1_core 91,976 92,827 1012 1066 974 1067 974 903 1290 873 624 631
neuron 92,290 125,305 252 260 244 262 243 206 270 244 270 244
stereo_vision 94,050 127,085 180 180 169 176 169 91 143 91 93 91
des90 111,221 139,557 402 402 380 372 379 358 441 380 349 345
SLAM_spheric 113,115 142,408 1061 1061 1061 1061 1061 1061 1061 1061 1061 1061
cholesky_mc 113,250 144,948 285 285 283 283 283 345 667 591 281 479
segmemtation 138,295 179,051 126 136 107 114 107 78 141 78 78 78
bitonic_mesh 192,064 235,328 585 614 593 594 586 483 590 592 493 506
dart 202,354 223,301 807 844 924 805 784 540 603 594 549 539
openCV 217,453 284,108 510 511 560 635 499 518 554 501 554 473
stap_qrd 240,240 290,123 399 399 371 386 371 295 295 275 287 275
minres 261,359 320,540 215 215 207 215 207 189 189 199 181 191
cholesky_bdti 266,422 342,688 1156 1157 1156 1161 1156 947 1024 1120 1024 848
denoise 275,638 356,848 416 722 416 516 416 224 478 244 224 220
sparcT2_core 300,109 302,663 1244 1273 1186 1319 1183 1245 1972 1186 1081 918
gsm_switch 493,260 507,821 1827 5974 1759 1714 1621 1407 5352 1719 1503 1407
mes_noc 547,544 577,664 634 699 649 699 651 617 633 755 633 617
LU230 574,372 669,477 3273 4070 4012 3452 3602 2677 3276 3751 2720 2923
LU_Network 635,456 726,999 525 550 524 550 524 524 528 524 528 524
sparcT1_chip2 820,886 821,274 899 1524 874 1129 873 783 1029 856 877 757
directrf 931,275 1,374,742 574 646 646 646 632 295 379 295 317 295
bitcoin_miner 1,089,284 1,448,151 1297 1570 1576 1562 1514 1225 1255 1287 1255 1282
V Conclusion
In this study, we introduced SHyPar, a multilevel hypergraph partitioning framework that enhances partitioning solutions and surpasses earlier studies in performance. We developed an innovative algorithm that incorporates spectral hypergraph coarsening techniques, leverages hyperedge effective resistances and flow-based community detection. Our comprehensive experimental analysis, conducted on real-world VLSI test cases, demonstrates that SHyPar consistently achieves a significant reduction in hypergraph partitioning cut size, improving results up to 15 percent compared to state-of-the-art methods.
VI Acknowledgments
This work is supported in part by the National Science Foundation under Grants CCF-2417619, CCF-2021309, CCF-2011412, CCF-2212370, and CCF-2205572.